OpenStudy (anonymous):

Describe the differences between the graph of y = –3(x + 7)2 – 10 and the standard position graph of y = x2.

OpenStudy (anonymous):

the first could be written as: -3y(x+7) - 10

OpenStudy (jjuden):

When you graph a parabola. the number outside of your () is opposite. so move your starting point to the left 7. then move down 10 (due to the negative 10) and your parabola will be opening down (down 3, over one) The difference would be that it is in the middle of graph, while the first is not

OpenStudy (anonymous):

http://www.drcruzan.com/Images/TransformationEquation.png look at this pic. your parent equation is y = x^2 then it is transformed into y = –3(x + 7)^2 – 10

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait nvm guy i remember how to go about solving this now but thanks

OpenStudy (anonymous):

y = x^2 (1) y = -3(x + 7)^2 - 10. (2) Differences between the 2 graph (1( and (2) a) The parabola (1) is upward (a > 0); the parabola (2) is downward (a < 0) b) Parabola (1) passes through the origin 0; parabola (2) passes through point (-7, -10) c) x-intercept of (1) is (0, 0); (2) doesn't have x-intercept d) Develop (2)-> y = -3(x^2 + 14x + 49) - 10 = -3x^2 - 42x - 157.y-intercept of (1) is (0, 0); y-intercept of (2) is (0, -157)